# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ 2.1 Differential Privacy
Let $X$ be the input dataset. Each data point $x_{i}$ is defined over some measurable domain $\mathcal{X}^{\dagger}$, so that $x=(x_{1},x_{2},\ldots,x_{n})\in\mathcal{X}=(\mathcal{X}^{\dagger})^{n}$. We say two datasets $x,x^{\prime}\in\mathcal{X}$ are adjacent if they differ in exactly one data point.
Definition 2.1 ($(\epsilon,\delta)$-Differential Privacy dwork2006calibrating).
@@
2.2 DP-SGD with Gradient Clipping
Differentially private stochastic gradient descent (DP-SGD) Abadi2016; song2013stochastic has become the standard approach for training neural networks with formal privacy guarantees.
@@ 2.2 DP-SGD with Gradient Clipping
Let $x=(x_{1},\dots,x_{n})\in\mathcal{X}=\bar{\mathcal{X}}^{n}$ be the finite samples of data points, where each $x_{i}=(a_{i},b_{i})\in\bar{\mathcal{X}}$ consists of features $a_{i}$ and label $b_{i}$. In this section, we take $P^{\theta}_{X}$ as the empirical (marginal) distribution of the data samples $x=(x_{1},\dots,x_{n})\in\mathcal{X}=\bar{\mathcal{X}}^{n}$ for all $\theta\in\Theta$. At iteration $t$, we draw a random mini-batch $\mathsf{I}_{t}\subseteq[n]$ using a subsampling scheme $\{\eta,\rho\}$ characterized by a rate $\eta$ and a scheme type $\rho$, where $\rho\in\{\textsf{WR},\textsf{WOR},\textsf{Poisson}\}$ denotes sampling with-replacement (WR), without-replacement (WOR), or via a Poisson process (Poisson), respectively. Let $X^{t}=(X_{i})_{i\in\mathsf{I}_{t}}$ denote the random mini-batch subsampled from $x$, with $x^{t}$ representing a particular realization.
@@
DP-SGD song2013stochastic; Abadi2016 modifies SGD in two ways: it clips per-example gradients to control sensitivity and adds noise to the averaged gradient for privacy. Each per-sample gradient is clipped at $\ell_{2}$ radius $C$,
@@
$\bar{g}_{t}:=\frac{1}{B}\sum_{i\in\mathsf{I}_{t}}\tilde{g}_{t}(x_{i}).$ (1)

### v3

@@ A.1 Differential Privacy
This section revisits the Differential Privacy (DP). Let $X$ be the input dataset. Each data point $x_{i}$ is defined over some measurable domain $\mathcal{X}^{\dagger}$, so that $x=(x_{1},x_{2},\ldots,x_{n})\in\mathcal{X}=(\mathcal{X}^{\dagger})^{n}$. We say two datasets $x,x^{\prime}\in\mathcal{X}$ are adjacent if they differ in exactly one data point.
Definition A.1 ($(\epsilon,\delta)$-Differential Privacy [dwork2006calibrating]).
@@
A.2 DP-SGD with Gradient Clipping
In this section, we describe the Differentially private stochastic gradient descent (DP-SGD) [Abadi2016, song2013stochastic] has become the standard approach for training neural networks with formal privacy guarantees. First, We provide the following basic setup for DP-SGD similar to Sec. 5 for completeness.
@@ A.2 DP-SGD with Gradient Clipping
Let $x=(x_{1},\dots,x_{n})\in\mathcal{X}=\bar{\mathcal{X}}^{n}$ be the finite samples of data points, where each $x_{i}=(a_{i},b_{i})\in\bar{\mathcal{X}}$ consists of features $a_{i}$ and label $b_{i}$. In this section, we take $P_{\theta}^{X}$ as the empirical (marginal) distribution of the data samples $x=(x_{1},\dots,x_{n})\in\mathcal{X}=\bar{\mathcal{X}}^{n}$ for all $\theta\in\Theta$. At iteration $t$, we draw a random mini-batch $\mathsf{I}_{t}\subseteq[n]$ using a subsampling scheme $\{\eta,\rho\}$ characterized by a rate $\eta$ and a scheme type $\rho$, where $\rho\in\{\textsf{WR},\textsf{WOR},\textsf{Poisson}\}$ denotes sampling with-replacement (WR), without-replacement (WOR), or via a Poisson process (Poisson), respectively. Let $X^{t}=(X_{i})_{i\in\mathsf{I}_{t}}$ denote the random mini-batch subsampled from $x$, with $x^{t}$ representing a particular realization.
@@
DP-SGD [song2013stochastic, Abadi2016] modifies SGD in two ways: it clips per-example gradients to control sensitivity and adds noise to the averaged gradient for privacy. Each per-sample gradient is clipped at $\ell_{2}$ radius $C$,
@@
$\bar{g}_{t}:=\frac{1}{B}\sum_{i\in\mathsf{I}_{t}}\tilde{g}_{t}(x_{i}).$